# RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions

### v1

@@ 4.2 Confusion Detection
+ self-consistency (n=3) 98.23
+ self-consistency (n=9) 98.23
Table 2: Evaluation of different prompts in the task “defusion detection” on RAG-ConfusionQA-Golden.
@@ 4.3 Ablation Study
We examined the impact of various prompts on the task of “defusion detection.” As seen in Table 2, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $n=3$ further boosted performance to 98.23%. However, increasing the number of generated sequences beyond three did not lead to additional improvements. Since RAG-ConfusionQA-Golden is a smaller dataset, using $n=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger RAG-ConfusionQA-Silver dataset. Therefore, we adopted $n=9$ in our experiments to balance accuracy and expenses (Appendix F.10).
@@ 5 Conclusion
We introduced RAG-ConfusionQA, a benchmark designed to assess how effectively an LLM defuses confusion when responding to confusing questions using a standard prompt. Through human annotation, we demonstrated that RAG-ConfusionQA achieves a 94.91% accuracy in generating confusing and non-confusing questions, while our proposed AutoDefusionEval achieves a 98.23% accuracy in detecting whether an LLM’s response defuses confusion. Additionally, our evaluation highlights that current LLMs often attempt to answer confusing questions to varying degrees. For future work, we aim to develop methods to better detect confusing questions and address the tendency of LLMs to respond to them directly instead of defusing the confusion.
@@ 7 Limitations
In this paper, we only evaluated LLMs’ abilities to generate responses that effectively defuse confusing questions when prompted with a standard prompt (Appendix F.6). We chose this prompt because, ideally, LLMs should be able to address confusion with a basic, straightforward prompt. However, we have not yet assessed the performance of LLMs across different prompt variations.
@@ Appendix A Algorithm
1:Call LLMr on $(d,q)$ with the standard RAG prompt and get its response $r$ (Appendix F.6)
2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is confusing given $d$ (Appendix F.3)
@@
4:if $q$ is generated as confusing then
@@ Appendix B Human Annotation
Figure 2: Confusion matrix of confusion and defusion on RAG-ConfusionQA-Golden.
@@
confusion defusion confusion defusion confusion defusion confusion defusion
@@
Table 3: Agreement scores and the performance of our method on the RAG-ConfusionQA-Golden.
Following the methodology from Cochran (1977), representing approximately 20,000 synthetic questions in RAG-ConfuseQA at a 95% confidence level, with a 4% margin of error, and an anticipated accuracy of $90\pm 5\%$, required labeling at least 202 questions. We sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators. As shown in Table 3, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is confusing and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency. Two more annotators manually resolved all disagreements to establish ground truth labels. Six annotators are computer science graduate students recruited via email and incentivized with complimentary meals. Another two annotators are the authors of this paper, who are also majoring in computer science.
@@ Appendix C Different LLMr
Table 4: Different LLMr to be evaluated. In RAG-ConfuseQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
Appendix D Confusing Types
@@
Question 1: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
This question is confusing as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document. So this belongs to Not Mention
Question 2: How will Justin Jefferson’s struggles with hamstring issues in Week 3 affect the Vikings’ chances of winning against the Houston Texans?
This question assumes that Justin Jefferson is dealing with hamstring issues, which is not mentioned in the document. The document only refers to a quad injury that Jefferson had, which is no longer a concern as of Week 3. The document makes it clear that Jefferson is healthy and will play without any restrictions. Therefore, the question is based on a False Premise
Question 3: How will his performance impact the team’s chances this season, considering they are missing key players?
This question uses the Ambiguous pronoun “his” without specifying who it refers to, making it unclear whether the question is asking about Justin Jefferson, CeeDee Lamb, or another player mentioned in the document. Additionally, “the team” is not defined explicitly, as both the Minnesota Vikings and other teams are referenced in the document. Without clarifying which player or team the question is referring to, it is impossible to provide an accurate answer based on the document.
Appendix E Statistics of RAG-ConfusionQA
Attr Golden Silver
# of documents 45 2,000
# of non-confusing questions 103 9,052
# of confusing questions 113 10,218
@@ Appendix E Statistics of RAG-ConfusionQA
average words per non-confusing question 16 16
average words per confusing question 16 16
Table 5: Statistics of RAG-ConfusionQA. Golden is sampled from Silver and annotated by annotators (Appendix B).
@@ F.1 Extract Claims Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and list {num_fact} most important facts it contains. Each fact should be
stated in a clear, standalone sentence with sufficient context to be understood
independently, avoiding undefined pronouns. Ensure that each fact is directly derived from
the document and does not include any information not mentioned within it.
@@
{num_fact} most important facts:"
@@ F.2 Recover Missing Claims Prompt
{mdframed}
Read the document below with a list of {num_fact} facts it contains. Note that some of the
facts are missing represented by (missing). Your task is to guess the missing facts could
have said and complete the missing facts. Each fact should be stated in a clear, standalone
sentence with sufficient context to be understood independently, avoiding undefined pronouns.
Please complete the missing facts and return all the {num_fact} facts in original order. You
must return all the {num_fact} facts.
@@
Completed list of facts:"
F.3 Confusion Judgement Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and the question. Check if the question contains any confusing
parts, false assumptions, incorrect information, or mentions of people, locations,
organizations, or other entities that are not present in the document.
@@ F.3 Confusion Judgement Prompt
Does the question contain any confusing parts, false assumptions, incorrect
information, or mentions of people, locations, organizations, or other entities
that are not present in the document? Please start by explaining your reasoning.
End your response with ’The answer is: Yes or No.’:
F.4 Defusion Judgement Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
@@ F.4 Defusion Judgement Prompt
    {example_question}
{defuse_answer_examples}
{Defusion Instruction}
{defuse_reasoning_examples}
@@
{response_from_gpt3.5}
{Defusion Instruction}
@@ F.5 Defusion Instruction
{mdframed}
The question presented is confusing and unanswerable based on the document provided.
Read the response carefully. If the response attempts to answer the question,
like answering ’Yes’ to agree the question or ’No’ to disagree the question,
you should immediately conclude with The answer is: No. even if it later
acknowledges the questions confusion. Otherwise, determine if the response
clarifies that the question cannot be answered or confusing or matches any of the
issues listed:
    False Premise: The response points out that the question is based
    on an incorrect assumption.
    Not Mentioned: The response indicates that the information needed to answer
    the question is not available or not mentioned in the document.
    Ambiguous: The response points out unclear pronouns or vague terms in the
    question that prevent a clear answer.
Begin with your reasoning and finally conclude with The answer is: Yes. or
The answer is: No.. Note that if the response attempts to answer the question
at any point, like answering ’Yes’ to agree the question or ’No’ to disagree the
question, you should immediately conclude with The answer is: No. even if it
later acknowledges the questions confusion:
F.6 RAG Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
@@ F.6 RAG Prompt
F.7 Non-confuing Question Generation Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document attentively and compile a numbered list of the top {num_q} questions
that the document directly answers. Ensure each question is clear, accurate, and devoid
of confusion, false assumptions, undefined pronouns, or misinformation. Avoid referencing
people, locations, organizations, or other entities not explicitly mentioned in the
document. Construct each question to be thought-provoking, containing between 13 to 18
words, and sufficiently detailed to avoid being overly straightforward.
@@ F.7 Non-confuing Question Generation Prompt
F.8 Confusing Question Generatioin Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and review the list of hallucinated facts. For each hallucinated fact,
craft a single, specific and concise question containing 13 to 18 words that incorporates
the key element of the fact, ensuring the question is intentionally confusing.
The question should not be answerable using any information present in the document.
The question should not combine multiple queries and each question should address only
one specific aspect. If a question cannot be formulated for a particular hallucinated
fact, you may omit it.
@@ F.8 Confusing Question Generatioin Prompt
F.9 Remove Claims Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document below with a list of {num_true_fact} ground-truth facts it contains
and a list of {num_false_fact} hallucinated facts that are not supported by the document.
Your task is to remove any hallucinated facts that can be supported by either the
document or the {num_true_fact} ground-truth facts.
Please only return the remaining hallucinated facts, along with their original order numbers.
@@ F.9 Remove Claims Prompt
4 ground-truth facts:
@@
Read the document below with a list of {num_true_fact} ground-truth facts it contains
and a list of {num_false_fact} hallucinated facts that are not supported by the document.
Your task is to remove any hallucinated facts that can be supported by either the
document or the {num_true_fact} ground-truth facts.
Please only return the remaining hallucinated facts, along with their original order numbers.
@@
{num_true_fact} ground-truth facts:
@@
F.10 Expenses
Generating the RAG-ConfusionQA dataset using GPT-4o-mini costs approximately \$50. For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/. Excluding the dataset generation, all other experiments cost around \$60.

### v2

@@ 4.2 Out-of-Scope Detection
+ self-consistency (m=3) 98.23
+ self-consistency (m=9) 98.23
Table 3: Evaluation of different prompts for “defusion detection” on ScopeQA-Gold.
@@ 4.3 Ablation Study
We examined the impact of various prompts on “defusion detection” task. As seen in Table 3, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $m=3$ further boosted performance to 98.23%. However, further increasing $m$ did not lead to additional improvements. Since ScopeQA-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ScopeQA-Silver dataset. Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix G).
@@ 5 Conclusion
We introduced ScopeQA, a benchmark for evaluating LLMs’ ability to defuse out-of-scope questions in RAG. Human annotation validates that ScopeQA achieves 94.91% accuracy in distinguishing out-of-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses. We further demonstrated ScopeQA’s utility by training out-of-scope detectors. For future work, we plan to investigate the impact of appending an unused-token on the generation process. Additionally, we aim to explore instruction fine-tuning by training LLMs on ScopeQA and evaluating their performance. We also intend to apply for OpenAI funding to leverage more powerful models, such as GPT-4o, to enhance data quality.
@@ 7 Limitations
In this paper, we evaluated LLMs’ abilities to defuse out-of-scope questions using three prompting methods: basic prompt, Two-shot prompt, and Zero-shot-CoT prompt. These methods were chosen because LLMs should ideally handle out-of-scope questions with straightforward prompts. However, we did not assess the performance of LLMs fine-tuned on ScopeQA-Silver due to computational resource constraints. Additionally, we empirically appended an unused-token to the end of the prompt for classification while minimizing its influence on LLMs’ generation processes. The effect of appending such tokens on generation remains unexplored and is left for future work. Finally, budget constraints prevented us from using more powerful models, such as GPT-4o or GPT-o1, which could theoretically produce higher-quality datasets. Future users can adopt our framework with more advanced models and alternative data sources to generate even higher-quality data tailored to their needs.
@@ Appendix A Algorithm
1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix F.6,  F.7,  F.8)
2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix F.3)
@@
4:if $q$ is generated as out-of-scope then
@@ Appendix B Human Annotation
Figure 2: Confusion matrix of out-of-scope and defusion on ScopeQA-Gold.
@@
Out-of-Scope defusion Out-of-Scope Defusion Out-of-Scope Defusion Out-of-Scope Defusion
@@
Table 4: Agreement scores and the performance of our method on the ScopeQA-Gold.
Following the methodology from Cochran (1977), we sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators. As shown in Table 4, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is out-of-scope and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency. Two more annotators manually resolved all disagreements to establish ground truth labels. All the annotators are computer science graduate students recruited via email and incentivized with complimentary meals. All annotators were trained by first reviewing the annotation guidelines and then labeling a small dataset previously labeled by the authors. The annotation guideline and annotation tool will be released along with the data and code upon acceptance.
@@ Appendix C Different LLMr
Table 5: Different LLMr to be evaluated. In ScopeQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
Appendix D Out-of-scope Example
@@
Question: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
Explanation: This question is out-of-scope as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document.
Appendix E Statistics of ScopeQA
Attr Gold Silver
# of documents 45 1955
# of in-scope questions 103 8,949
# of out-of-scope questions 113 10,105
@@ Appendix E Statistics of ScopeQA
average words per in-scope question 16 16
average words per out-of-scope question 16 16
Table 6: Statistics of ScopeQA. Gold is sampled from crawled 2000 news and annotated by annotators (Appendix B).
@@ F.1 Extract Claims Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and list {num_fact} most important facts it contains. Each fact should be stated in a clear, standalone sentence with sufficient context to be understood independently, avoiding undefined pronouns. Ensure that each fact is directly derived from the document and does not include any information not mentioned within it.
@@
{num_fact} most important facts:
@@ F.2 Recover Missing Claims Prompt
Read the document below with a list of {num_fact} facts it contains. Note that some of the facts are missing represented by (missing). Your task is to guess the missing facts could have said and complete the missing facts. Each fact should be stated in a clear, standalone sentence with sufficient context to be understood independently, avoiding undefined pronouns. Please complete the missing facts and return all the {num_fact} facts in original order. You must return all the {num_fact} facts.
@@
Completed list of facts:
F.3 Out-of-scope Judgement Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and the question. Check if the question contains any confusing parts: mentions of people, locations, organizations, or other entities that are not present in the document
@@ F.3 Out-of-scope Judgement Prompt
Does the question contain any confusing parts: mentions of people, locations, organizations, or other entities that are not present in the document? Please start by explaining your reasoning. End your response with ’The answer is: Yes or No.’:
F.4 Response Defusion Judgement Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document and answer the question based on the document.
@@ F.4 Response Defusion Judgement Prompt
{example_questions}
Answer:
{example_LLM_responses}
{Defusion Instruction (Appendix  F.5)}
{example_defuse_reasoning}
@@
Answer:
{LLM_response}
{Defusion Instruction (Appendix  F.5)}
@@ F.5 Defusion Instruction
The question presented is confusing and unanswerable based on the document provided. Read the response carefully. If the response attempts to answer the question, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion. Otherwise, determine if the response clarifies that the question cannot be answered or if the response indicates that the information needed to answer the question is not available or not mentioned in the document. Begin with your reasoning and finally conclude with ‘The answer is: Yes.’ or ‘The answer is: No.’. Note that if the response attempts to answer the question at any point, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion:
F.6 RAG Basic
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@ F.6 RAG Basic
Answer:
F.7 RAG Two-shot
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
"""{example_document}"""
Question:
{two_example_questions}
Answer:
{two_example_responses}
Read the document and answer the question based on the document.
Document:
"""{document}"""
Question:
{question}
Answer:
F.8 RAG Zero-shot-CoT
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and reason step by step to answer the question based on the document. If the question cannot be answered using the document, state explicitly that the question cannot be answered.
Document:
"""{document}"""
Question:
{question}
Answer:
F.9 In-scope Question Generation Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document attentively and compile a numbered list of the top {num_q} questions that the document directly answers. Ensure each question is clear, accurate, and devoid of confusion, false assumptions, undefined pronouns, or misinformation. Avoid referencing people, locations, organizations, or other entities not explicitly mentioned in the document. Construct each question to be thought-provoking, containing between 13 to 18 words, and sufficiently detailed to avoid being overly straightforward.
@@ F.9 In-scope Question Generation Prompt
F.10 Out-of-scope Question Generation Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document and review the list of hallucinated facts. For each hallucinated fact, craft a single, specific and concise question containing 13 to 18 words that incorporate the key element of the fact, ensuring the question is intentionally confusing. The question should not be answerable using any information present in the document. The question should not combine multiple queries and each question should address only one specific aspect. If a question cannot be formulated for a particular hallucinated fact, you may omit it.
@@ F.10 Out-of-scope Question Generation Prompt
F.11 Remove Claims Prompt
You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
Read the document below with a list of {num_true_fact} ground-truth facts it contains and a list of {num_false_fact} hallucinated facts that are not supported by the document. Your task is to remove any hallucinated facts that can be supported by either the document or the {num_true_fact} ground-truth facts. Please only return the remaining hallucinated facts, along with their original order numbers.
@@ F.11 Remove Claims Prompt
{num_true_fact} ground-truth facts:
@@
Read the document below with a list of {num_true_fact} ground-truth facts it contains and a list of {num_false_fact} hallucinated facts that are not supported by the document. Your task is to remove any hallucinated facts that can be supported by either the document or the {num_true_fact} ground-truth facts. Please only return the remaining hallucinated facts, along with their original order numbers.
@@
{num_true_fact} ground-truth facts
@@
Appendix G Expenses
Generating the ScopeQA dataset using GPT-4o-mini costs approximately \$80. For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/. Excluding the dataset generation, all other experiments cost around \$100.
Appendix H Embedding Classifier
H.1 Implementation Details
We train the classifier on ScopeQA-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ScopeQA-Gold data.
We append the unused-token ID at the end of the tokenized input sequence. After the LLM processes the input, we extract the unused-token’s hidden state from the model’s last hidden layer output. Since this token is not part of the LLM’s training and does not carry predefined meanings, it serves as a placeholder to aggregate information from the model’s attention over all preceding tokens in the sequence. As the final token in the input, its embedding encapsulates the model’s representation of the entire input sequence, reflecting the information inferred from the input sequence.
The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss. A dropout of 0.1 is applied after the first layer. We use Adam optimizer with 1e-4 learning rate. We train each classifier for 10 epochs with batch size 8. We save the best validation checkpoint and report its performance on ScopeQA-Gold.
H.2 Unused-Token
For the Llama family, we use the <|reserved_special_token_0|> as the unused-token. This token is not trained and is reserved for future use cases such that the users do not need to resize the vocabulary. For Mistral-7B-v0.3, we use [control_555] as the unused-token because Mistral’s official documentation indicated that they reserve 768 control tokens for future use, and none of them are used during training. 555 is a random choice.
H.3 Sentence BERT NLI baseline
To create a baseline classification performance, we adapt NLI style training using an off-the-shelf sentence-transformers/all-mpnet-base-v2 model from HuggingFace. Specifically, we follow Reimers and Gurevych (2019) to extract document and question [CLS] embeddings $u$ and $v$ to use $[u;v;|u-v|]$ as the input feature for the 2-layer MLP to classify.